### v1

# Multiple References with Meaningful Variations Improve Literary Machine Translation

Abstract
While a source sentence can be translated in many ways, most machine translation (MT) models are trained with only a single reference. Previous work has shown that using synthetic paraphrases can improve MT. This paper investigates best practices for employing multiple references by analyzing the semantic similarity among different English translations of world literature in the Par3 dataset (Thai et al., 2022). We classify the semantic similarity between paraphrases into three groups: low, medium, and high, and fine-tune two different LLMs (mT5-large and LLaMA-2-7B) for downstream MT tasks. Across different models, holding the total training instances constant, single-reference but more source texts only marginally outperforms multiple-reference with half of the source texts. Moreover, using paraphrases of medium and high semantic similarity outperforms an unfiltered dataset (+BLEU 0.3-0.5, +COMET 0.2-0.9, +chrF++ 0.25-0.32). Our code is publicly available on GitHub. [^1] https://github.com/swsiwu/multi_ref_literary_MT
Multiple References with Meaningful Variations Improve
Literary Machine Translation
Si Wu Northeastern University siwu@ccs.neu.edu                        John Wieting Google DeepMind jwieting@google.com                        David A. Smith Northeastern University dasmith@ccs.neu.edu
1 Introduction
A sentence can be translated into another language in many ways. The differences among translations can be trivial—replacement of a synonym, punctuation, possessive ’s vs. of—or they can be meaningful and reveal different aspects of the source language text. (See examples in Figure 1, Table 5, and Appendix B.)
Literary translators have to consider multiple aspects of a source language when translating: meaning, form, rhyme, style, corresponding culture that the language is embedded in, etc. (Lefevere, 1992). Some emphasize lexical and syntactic imitation (formal equivalence), while others sacrifice form for comprehensibility (dynamic/functional equivalence) (Nida, 1964). Read together, different translations’ emphases and interpretations provide a more comprehensive impression of the source text and a deeper understanding of the source language.
Figure 1: An example of meaningful variations between different English translations of a Classical Chinese source text. Highlighted texts are segments where the variations are meaningful. The example source text and its English translations are from the Par3 dataset (Thai et al., 2022).
Previous work has shown that using synthetic paraphrases can improve MT performance (Madnani et al., 2008; Zheng et al., 2018; Khayrallah et al., 2020), but they do not explore different semantic similarity of paraphrases and their impact on MT performance.
To systematically investigate this, we use the Par3 dataset, which has paragraph-level, complex English paraphrases from literature translated by expert translators (Thai et al., 2022). While there are highly similar paraphrases in Par3, many variations between paraphrases in Par3 are more than replacement, insertion, deletion, and re-ordering of words. These naturally existing paraphrases are more complex than what’s generated by traditional paraphrasers that are rule-based (McKeown, 1983) and SMT-based (Wubben et al., 2010; Bannard and Callison-Burch, 2005; Madnani et al., 2007). While LLMs have improved synthetic paraphrase’s quality and complexity, they are still at risk of mistakes, especially for outputs longer than one sentence (Jayawardena and Yapa, 2024a, b; Su et al., 2022; Horvitz et al., 2024; Mishra et al., 2024). Moreover, non-expert translations are prone to translationese, where the translation is overly literal and awkward (NLLB Team et al., 2022).
In this paper, we investigate two questions: when holding the amount of fine-tuning data constant, is using multiple references better than just one target reference? What kind of paraphrase should be incorporated to achieve the best performance?
To answer these questions, we create datasets that vary in the number of source texts, the number of references per source text (single vs multiple references), and for multiple-reference datasets, vary in the level of semantic similarity among the references. We use Paragram-SP (Wieting et al., 2022), an efficient and effective model for measuring semantic similarity. We divide paraphrases into three groups based on their semantic similarity: low, medium, and high, where medium has the most interesting and meaningful variations among the references for a source text. We then fine-tune two LLMs, mT5-large and LLaMA-2-7B, using aforementioned datasets on MT tasks, and compare their performances with baseline and commercial models: No Language Left Behind (NLLB Team et al., 2022) and Google Translate.
Our findings in the literary domain are:
  • Holding the total training instances constant, single-reference but more source texts only marginally outperforms multiple-reference. (Section 6.1)
  • Using source texts with medium and high semantic similarity among references can outperform an unfiltered dataset even if the unfiltered dataset has more training instances. (Section 6.2)
  • On the other hand, training with medium and low semantic similarity among references will perform worse than an unfiltered dataset. (Section 6.2)
The remainder of our paper is organized as follows: we describe our method for measuring semantic similarity in Section 3, then we describe how we construct different datasets based on their semantic similarity and number of source and reference texts in Section 4. We provide details on the models we use for fine-tuning as well as the baseline MT models in Section 5, where we also provide information about the MT metrics we use and relevant training details. Finally, we discuss our experimental results in Section 6.
2 Related work
Paraphrases have been used to augment training data and improve MT performance. Madnani et al. (2008) explore how much improvement one can get by increasing the number of synthetic references when tuning an SMT system. The paraphrases they use are automatically generated using phrase table pivoting. Similarly, Callison-Burch et al. (2006) also use pivot-based method to generate paraphrases to improve SMT performance. However, neither of these works examines the linguistic variations between the reference translations of a source text, which is reasonable since the references are synthetic and much less complex.
Khayrallah et al. (2020) use simulated paraphrases to improve MT performance on low-resource languages. They focus on the general effect of augmenting training data using synthetic paraphrases when training data is scarce. Our work further studies the semantic similarity of paraphrases on MT. It is worth noting that the languages present in our dataset are not low-resource languages, and instead of training a language model from scratch, we fine-tune two different types of LLMs: mT5 (an encoder-decoder model) and LLaMA-2-7B (an autoregressive LLM).
Zheng et al. (2018) propose a method to compress existing human references into a lattice and then generate synthetic references for training to boost MT and text generation performance. They also explore different ways of utilizing multiple references during training with a bidirectional language model. Instead of exploring methods of creating synthetic references, our work uses all available expert references for each source text.
Many studies have operationalized paraphrase taxonomy with different levels of granularity (Bhagat and Hovy, 2013; Wahle et al., 2023; Wegmann et al., 2024). Most are trained on human-annotated paraphrase datasets with explicit definitions and follow different agreement protocols among human annotators. Wegmann et al. (2024) operationalize context-dependent paraphrases with plausible variations in dialog, and they demonstrate their hands-on training approach for annotators. With this annotated dataset, they use in-context learning to train a classifier to detect paraphrases in dialog. In our paper, instead of using human annotations, we use Paragram-SP, a semantic textual similarity scorer, to classify the type of paraphrase, as we are interested in a computational, low-cost, yet effective approach.
Others explore different factors that can impact translation quality: Khayrallah and Koehn (2018) explore how different kinds of noise in training data impact MT performance when using an NMT model. Maillard et al. (2023) show that the addition of high-quality data can significantly improve MT performance on low-resource languages, even when the quantity of high-quality training data is small. Consequently, they emphasize the value of collecting high-quality data from native speakers for training MT models.
3 Measuring semantic similarity
We want to investigate what level of semantic similarity among the references (of a source text) is best for training MT systems by first measuring and classifying it.
Given a source text, we use a semantic textual similarity scoring model, Paragram-SP (Wieting et al., 2022), to measure the semantic similarity between its English translations. Paragram-SP score ranges from -1 to 1, where a lower score indicates low semantic similarity, and a higher score indicates high semantic similarity. A model similar to this one from Wieting et al. (2019b) was successfully used as a reward function in minimum risk training Shen et al. (2016) to improve machine translation quality Wieting et al. (2019a).
In the literary translation dataset Par3 (Thai et al., 2022), each source text has at least two different English translations. Paragram-SP can only score the semantic textual similarity for two texts, so in the case where there are more than two translations, we take the average pairwise Paragram-SP score as the source text’s sim_p score. Mathematically:
[…]
where $M$ is the set of human translations given a source text $s$, and $C$ is all possible combinations of a set of pairs in $M$. Paragram-SP (a,b) is the Paragram-SP score given a pair of English translations $a$ and $b$.
Figure 2: The distribution of sim_p for the Par3 corpus’s source texts. Paraphrases with high scores are more similar semantically and syntactically. Upon manual inspection, we observe that the paraphrases in between the two blue vertical lines are the ones with meaningful variations: different enough to be interesting, but similar enough to still be a paraphrase.
To identify the range for meaningful variations, we rank the source texts by their sim_p scores from low to high with their target references attached. Upon manual inspection, we define the range for meaningful variations to be [0.45-0.85], and we call this set of paraphrases the Medium dataset. Subsequently, the lower range [-1,0.45] is the Low dataset, and [0.85,1] is the High dataset.
We provide a detailed description of each sim_p range in Table 1, and the distribution of sim_p in the entire unfiltered dataset is in Figure 2.
sim_p Description Dataset
$1.0$ Completely identical. Medium+High
Unfiltered
$[0.9-1.0)$ Almost identical, except for minor punctuation differences, Medium+High
e.g. single vs double quotation marks, or word replacement Unfiltered
(by a synonym or a phrase).
$[0.85-0.9)$ Differences happen more often, but they are still variations Medium+High
that are not meaningful. Compared to the 0.9-1.0 similarity, Unfiltered
the difference can be clausal instead of just one or two words.
However, the overall amount of information is about the same.
$[0.45-0.85)$ Very meaningful variations. The paraphrases are more likely Medium
to be visually dissimilar. Sometimes, the amount of information Medium+High
can differ much more: some translators are more verbose, Medium+Low
others are more succinct. Unfiltered
$[-1-0.45)$ Noisy paraphrases that have an overall information mismatch. Medium+Low
Paraphrases can be misaligned or erroneous. Unfiltered
Table 1: Different levels of semantic similarity score (sim_p) and their characteristics. See Section 3 for how we compute sim_p given a source text and its reference translations. The dataset column indicates the datasets that include paraphrases from the corresponding sim_p range.
4 Datasets
We classify reference translations based on their semantic similarity with the method described in Section 3, then we create different datasets composed of various sim_p for experiments.
To systematically study the impact of the number of source texts, the number of references, and the type of paraphrase included, we create different training datasets such that each varies one potentially influential factor.
We use #source to denote the number of unique source texts, and #instance to denote the total number of training instances, e.g. 3 source texts each of which has 3 reference translations, then #source = 3, and #instance = 9. #source and #instance for each dataset are in Table 2. Here we describe some representative datasets:
Medium All
#source = 49,650: This dataset uses multiple references for each source text. It contains all the paraphrases with sim_p $\in[0.45,0.85]$, of which we define to have the most meaningful variations (§3).
Medium
#source = 20,897: This dataset uses multiple references for each source text and has the same amount of training instances (#instance) as Single Ref #source = 49,650, but its #source is smaller. This dataset is created to find out the extent of the efficiency of training with Medium datasets. Similarly, we create Medium with #instance = 57,987 and 76,938 to compare with the corresponding single-reference datasets.
Unfiltered
#source = 49,650: This dataset uses multiple references for each source text, and it contains the same amount of source texts as Medium All, but its source texts are randomly selected regardless of their corresponding target texts’ sim_p.
Single Ref
#source = 49,650: This dataset has the same amount of source texts as Medium All, but each source text is only trained with one reference.
Medium+Low 10%:
This dataset uses multiple references for each source text, and it includes the entire Medium All dataset. Additionally, we randomly select 10% of the source texts from the lower sim_p range ($[0,0.45]$).
Medium+Low All:
Similar to Medium+Low 10%, but it includes all the source texts with sim_p in the lower range ($[0,0.45]$).
Medium+High 10%:
Similar to Medium+Low 10%, except the additional 10% source texts are from high sim_p range ($[0.85,1]$). The Medium+High datasets are to examine the impact of including less meaningful variations (high semantic similarity) when fine-tuning an MT model.
The source paragraphs and human references are from Par3 (Thai et al., 2022). [^2] Par3 dataset can be downloaded from https://github.com/katherinethai/par3 The source languages include Portuguese, Swedish, Italian, Czech, Japanese, Classical Chinese, Norwegian, German, Russian, and French. All references are only in English; In other words, the language pairs in this dataset are all to-English language pairs. We provide details about the composition of Medium all compared to Unfiltered All in Appendix 4.
We use paragraphs as units for source and target and only select source paragraphs whose human translations are all under the 128-word limit. Voigt and Jurafsky (2012) have shown that when applying MT to literary translation, using above sentence-level translation can improve literary cohesion. Moreover, a paragraph provides enough length for interesting variations but also increases the complexity of the task.
To ensure train/val/test has the same language split, we split data into train and val/test in 0.8/0.2 ratio for each language by the number of source paragraphs instead of the source-target pairs. Train and val/test also don’t share the same books. Val and test datasets are created by first randomizing at the source paragraph level. The test dataset is used in all experiments, and it has multiple human references per source text for evaluation. It is composed of medium semantic similarity only.
5 Experiments
We are interested in the effect of varying the number of source text, the number of references, and the semantic similarity of reference translations on MT tasks in the literary domain. Using datasets described in Section 4, we fine-tune two different general-purpose LLMs for translation: mT5-large and LLaMA-2-7B.
To calibrate our interpretation of the performance of these models, we compare them to out-of-the-box translation results from No Language Left Behind (NLLB) and Google Translate (§5.1). We employ three different translation evaluation metrics (§5.2). In the remainder of this section, we describe our training protocols (§5.3).
5.1 Models
mT5-large
(Xue et al., 2021): an encoder-decoder model that covers 101 languages (Xue et al., 2021) including all the languages in Par3. It is not specifically trained on translation tasks but could be fine-tuned for a variety of downstream tasks including multilingual translation. It has about 1.2 billion parameters. We choose to fine-tune mT5-large for our translation tasks. [^3] https://huggingface.co/google/mt5-large
LLaMA-2-7B
(Touvron et al., 2023): an English-centric LLM with limited multilingual capabilities. Studies have shown that LLaMA-2 can translate high-resource non-English language with some fine-tuning due to the presence of the corresponding language in the training dataset. LLaMA-2 is also able to adapt to a new low-resource language with some fine-tuning. All the languages in this paper are in the LLaMA-2 training data for >= 0.05%, except that the Chinese in Par3 is classical Chinese, which shares mostly the same characters as modern Chinese but has different grammar. It would be interesting to observe the translation performance after fine-tuning on our datasets given that non-English data is only a minuscule part of LLaMA-2 training data.
No Language Left Behind (NLLB):
An open-source MT system that can translate any language pairs between 202 languages, including low-resource languages, such as Asturian, Luganda, Urdu (NLLB Team et al., 2022). All the NLLB models are able to translate out-of-the-box, unlike mT5 which requires fine-tuning for MT downstream tasks. To compare to a fine-tuned mT5-large (1.2B), we choose the NLLB-200-Distilled model with 1.3B parameters. [^4] https://huggingface.co/facebook/nllb-200-distilled-1.3B We choose NLLB as a baseline model to compare to.
Google Translate:
A closed-sourced commercial system. As far as we know, it is trained on data from the public web with a transformer encoder and an RNN decoder. [^5] https://research.google/blog/recent-advances-in-google-translate/ Google Translate’s output is provided in the Par3 dataset.
5.2 Evaluation metrics
BLEU
(Papineni et al., 2002): An n-gram-based MT evaluation metric. Notice that we are using multiple human references for each source input.
$\text{COMET}_{22}^{\text{DA}}$
(Rei et al., 2020): An MT evaluation metric that is based on sentence embeddings. We use wmt22-comet-da by Unbabel (Rei et al., 2022). [^6] https://huggingface.co/Unbabel/wmt22-comet-da It was trained on a transformer-based multilingual masked language model, XLM-R (Conneau et al., 2020). Since $\text{COMET}_{22}^{\text{DA}}$ only accepts one reference per source text for evaluation, for each source text, we take the average of their $\text{COMET}_{22}^{\text{DA}}$ scores regarding different references translations.
chrF++
(Popović, 2017): An MT evaluation metric that is based on character and word-level n-gram recall and precision. We include this metric as Zouhar et al. (2024) find that COMET can struggle in domains outside WMT, e.g. literary domain.
#source #instance BLEU $\text{COMET}_{22}^{\text{DA}}$ chrF++
NLLB-1.3B 27.3 0.689 41.62
Google Translate 35.8 0.726 48.02
mT5 LLaMA-2 mT5 LLaMA-2 mT5 LLaMA-2
medium 20,897 49,650 29.8 30.0 71.3 72.8 43.20 44.06
single ref 49,650 49,650 29.9 30.1 71.5 70.3 43.44 42.08
unfiltered 49,650 116,689 31.3 30.8 71.8 72.7 44.17 44.52
medium all 49,650 118,003 31.5 30.3 71.9 71.6 43.81 43.76
medium 24,398 57,987 29.8 30.2 71.3 72.4 43.36 43.79
single ref 57,987 57,987 30.4 30.2 71.5 72.1 43.81 42.86
medium 32,357 76,938 30.6 28.2 71.6 70.2 43.73 41.12
single ref all 76,938 76,938 31.3 29.4 71.8 71.3 44.33 42.81
unfiltered 57,987 136,356 31.7 30.4 71.9 72.2 44.63 44.13
medium + low (10%) 57,987 137,811 31.1 29.4 71.7 71.6 44.16 41.29
medium + low (all) 59,072 140,407 31.1 30.1 71.7 71.7 44.11 43.64
unfiltered all 76,938 180,927 31.9 30.8 71.9 72.1 44.67 44.11
medium + high (10%) 57,987 136,870 32.0 29.2 72.0 70.3 44.77 42.39
medium + high (20%) 66,324 155,809 32.2 30.8 72.1 72.6 44.92 44.43
medium + high (all) 67,516 158,523 32.4 31.1 72.1 73.0 44.84 44.31
Table 2: Different datasets’ performance across different models. #source is the number of source texts, and #instance is the number of total training instances. For each MT metric, $\dagger$ denotes statistical significance for experiments using mT5, and $\ddagger$ denotes statistical significance for experiments using LLaMA-2-7B. We tested the statistical significance (p-value < 0.05) using the paired bootstrap resampling method (Koehn, 2004).
5.3 Training details
Given many LLMs already have strong language abilities and only require fine-tuning for MT downstream tasks, we consider fine-tuning as a more efficient approach than training from scratch, especially since our dataset is English-centric (all reference translations are in English) and these LLMs usually perform better when translating from non-English language into English.
For fine-tuning, each training instance is a source text and one of its human translations, and he training and validation datasets are randomized, e.g. (source text A, target human translation $A_{1}$), (source text B, target human translation $B_{3}$), (source text C, target human translation $C_{1}$), etc.
We fine-tuned mT5-large with Hugging Face libraries. Our optimal hyperparameters for fine-tuning are $lr=0.00025$, global batch size $=128$, and batch size $=16$. Training took 16 to 48 hours on 4 GPUs depending on the size of the dataset.
For LLaMA-2-7B, we used LoRA-tuning (Hu et al., 2022) with Meta’s LLaMA-recipe. [^7] https://llama.meta.com/llama-downloads/, [^8] https://github.com/meta-llama/llama-recipes We structured each training instance as "source_text ###> target text" following the fine-tuning method from Zhang et al. (2023). We used a batch size of 16. Depending on the dataset size, training took 8 to 24 hours.
For both models, we use BLEU on the validation set for early stopping. We notice that unlike mT5, the LLaMA-2 fine-tuning hyperparameters impact the performance significantly (±$\pm 5$ BLEU score). We also discover that when the training data increases significantly, increasing global batch size seems to improve the performance. In this paper, we report the best performance we obtained with three different global batch sizes for LLaMA-2-7B.
6 Results and Discussion
Using the datasets, models, and MT metrics described in previous sections, we analyze the effectiveness of the number and semantic similarity of reference translations.
Figure 3: Performance difference between different languages under different MT metrics when fine-tuning with mT5-large (1st row) and LLaMA-2-7B (2nd row). Darker dots are from Medium+High All, and lighter dots are from Unfiltered All. The bar chart show the amount of source-target-text pairs in Unfiltered All dataset.
6.1 Single reference vs multiple references
Simply adding additional references to a single-reference training set will, unsurprisingly, improve performance, just as will happen with increases in same-quality training data (Madnani et al., 2007, 2008). The more interesting question, however, is when the number of training instances are the same, does it make a difference whether we increase the amount of source text and pair each with one translation, or we have less source text but each source text is paired with multiple translations?
This is not, to our knowledge, an idea that has been previously explored. In Table 2, when n_instance is the same, Single Ref on average only marginally outperforms Medium across different models and metrics: ranging 0.0-0.7 in BLEU, (-)2.5-(+)1.1 in COMET, and (-)1.98-(+)1.69 in chrF++. It is surprising considering Medium has less than half of Single Ref’s n_source.
This suggests that when one has limited source and target texts, augmenting the number of references per source text can improve performance similarly to increasing the number of source text single target text pairs. However, it is unknown to what extent this observation will still hold when the references are synthetically generated with paraphrasers instead of high-quality human references in the literary domain.
6.2 Unfiltered All vs. Medium+High All
Although Unfiltered All has more training instances, its MT performance across different models and metrics is worse than Medium+High All and Medium+High 20% (Table 2).
We perform the paired bootstrap resampling to test the statistical significance of our finding (Koehn, 2004) and find that Med+High All outperforms Unfiltered All across different MT metrics and LLMs (p-value < 0.05). Specifically, Medium+High All outperforms Unfiltered All by 0.3-0.5 BLEU, 0.2-0.9 COMET, and 0.17-0.2 chrF++ (Medium+High 20% by 0.25-0.32 chrF++). This implies that when fine-tuning an MT model with multiple references, it is beneficial to compute the sim_p and filter out the ones with low semantic similarity to achieve the optimal MT performance.
Furthermore, we observe that Medium+Low 10% and All perform worse than Unfiltered All across different models and metrics (Table 2). Too much divergence among references of a source text may create confusion and hinder the performance.
6.3 Performance difference between models
We mentioned in Section 5.3 that LLaMA-2-7B is sensitive to the batch size when fine-tuning, and overall the fine-tuning result is less stable than mT5. In Table 2, we notice that BLEU, COMET, and chrF++ scores all seem less stable under LLaMA-2-7B, e.g. for Medium, increasing #instance resulted in a decrease in these metrics when #instance went from 57,987 to 76,938.
BLEU $\text{COMET}_{22}^{\text{DA}}$ chrF++
mT5 LLaMA-2-7B mT5 LLaMA-2-7B mT5 LLaMA-2-7B
medium + 1.0 31.7 28.9 72.0 71.3 44.59 42.46
medium + 0.9-1.0 31.7 27.8 71.9 70.3 44.55 40.75
medium + 0.85-0.9 31.7 31.0 71.9 72.9 44.52 44.70
Table 3: Ablation study on fine-tuning with the highly similar paraphrases (§6.5). 1.0 is pure repetition, 0.9-1.0 is the paraphrases with trivial differences, and 0.85-0.9 is non-trivial but non-meaningful paraphrases. All experiments have the same amount of source texts, #source = 57,987. For each MT metric, $\dagger$ denotes statistical significance for experiments using mT5, and $\ddagger$ denotes statistical significance for experiments using LLaMA-2-7B. We tested the statistical significance (p-value < 0.05) using the paired bootstrap resampling method (Koehn, 2004).
LLaMA-2’s MT ability is still not well understood: Xu et al. (2024) observe that for LLaMA-2-7B, increasing training data, under COMET, the MT performance declines, while under BLEU, it improves. From their observation, they suggested that for LLaMA-2-7B, increasing parallel data might dilute its pre-existing knowledge.
Although it is unknown what caused such unstable performance for LLaMA-2-7B, overall our observation in Section 6.2 still applies: using Medium+High is better than the unfiltered dataset, despite unfiltered dataset has larger #instance.
6.4 Performance difference between languages
All source languages in our datasets are also in the mT5-large and LLaMA-2-7B training data. However, as described in Section 5.1, non-English languages are only of an extremely small percentage of the total training data.
We compare the gain between Unfiltered All and Medium+High All among different languages under different metrics in Figure 3, and we also compare this gain with respect to the amount of fine-tuning data in Unfiltered All.
While among different metrics, Medium+High All outperforms Unfiltered All (Table 2), under the same model, each language’s metric gain is different. Surprisingly, for both mT5-large and LLaMA-2-7B, more fine-tuning data for a language (e.g. French and Russian have more fine-tuning data) does not result in proportionally more gain.
The gain for languages with less fine-tuning data also seems inconsistent between mT5-large and LLaMA-2-7B in both the amount and direction, e.g. Classical Chinese mostly has negative gain for mT5-large while much bigger positive gain using LLaMA-2-7B. However, overall, between metrics, the direction of gain for a language under the same model, is consistent.
6.5 Ablation study on the highly similar paraphrases
The Medium+High datasets (Medium+High 10%, 20% and ALL) are only different by the quantity of high semantic similarity paraphrases (Table 2 and Section 6.2,), in this ablation study, we further examine the characteristics of these high semantic similarity paraphrases and their impact on MT performance. We classify these high semantic similarity paraphrases into 3 categories:
  • sim_p = 1.0: completely identical paraphrases (absolute repetition)
  • sim_p [0.91.0$\in$[0.9-1.0)∈ [ 0.9 - 1.0 ): little variation, mostly lexical: 1 or 2 word replacements, deletion, addition, reordering, and punctuation change.
  • sim_p [0.850.9$\in$[0.85-0.9)∈ [ 0.85 - 0.9 ): slightly less trivial variations than the above. More syntactic variations, but still not meaningful.
For each category above, we combine Medium All and 10% additional paraphrases from that category and create a new dataset. These new datasets have the same number of source texts, #source = 57,987), with the exception that the first category (sim_p = 1.0) does not have enough data. To counter that, since the paraphrases in the first category are pure repetition, we create more training instances by randomly selecting source texts from the other two categories and repeating one of its reference translations.
Across different metrics, mT5 fine-tuned on Medium+1.0 provides the best performance, while LLaMA-2-7B performs best with Medium+0.85-0.9 (Table 3). We tested the statistical significance (p-value < 0.05) of the results in Table 3 using the paired bootstrap resampling method (Koehn, 2004).
Given the conflicting results from the two LLMs in this ablation study, it is inconclusive what the best kind of high semantic similarity is for achieving the best MT performance. We recommend simply including all medium and high semantic similarity paraphrases (Medium+High All) as suggested in Section 6.2.
7 Conclusion
Unlike more utilitarian domains of translation, literary works encourages multiple translators to try their hand at improving on their predecessors. To make the most of this valuable resource for training MT systems, we explore the optimal way to employ multiple references when fine-tuning LLMs for machine translation.
We find using source texts with medium and high semantic similarity among their references outperform an unfiltered dataset and provide the best performance.
Additionally, we observe that when holding the total number of training instances constant, single-reference (having more source texts with one target text each) is slightly better than multiple-reference (less source text with multiple target texts each). Future research could test these findings in other MT domains.
8 Limitations
  1. 1.
    We chose to fine-tune LLMs as a more efficient approach, given that these LLMs already have strong language ability and only require fine-tuning for downstream tasks. It is unclear if our claims still hold when one trains an LLM from scratch. Informed by the finding by Zhang et al. (2023), we did not use few-shot and zero-shot methods with LLaMA-2 because LoRA-tuning provides the best MT performance, and we considered LoRA-tuning to be the most appropriate to compare with fine-tuning the mT5-large model. Future research can study the performance difference among these approaches.
  2. 2.
    Fine-tuning mT5-large and LLaMA-2-7B usually requires multiple GPUs. Specifically, we trained mT5-large with 4 A100 GPUs since the model is very large and requires at least 30GB RAM. This could make it difficult for other researchers to replicate our experiments.
9 Potential Risks
MT systems are prone to gender-biased translation errors (Stanovsky et al., 2019), which could happen at test time in this work. Our dataset is also very English-centric since Par3 dataset only contains English paraphrases as target texts. English as a high-resource language is more likely to have multiple translations of the same book than low-resource languages, so working on English only will further widen the performance gap between high and low-resource languages; however, we hope our findings on the characteristics of paraphrases can be used in future research to improve MT on low-resource languages.
Appendix A Medium All dataset statistics
Lang book #src total %used
Portuguese 4 1021 3261 0.313
Swedish 3 1249 2591 0.482
Italian 2 1325 1912 0.693
Czech 4 1759 2886 0.609
Japanese 9 1779 2483 0.716
Chinese 7 2199 3671 0.599
Norwegian 2 3011 4756 0.633
German 16 6136 8905 0.689
Russian 27 20003 34030 0.588
French 32 32353 51052 0.634
Table 4: Composition of the Medium All dataset. "#src" is the total number of unique source paragraphs selected, while column "total" is the all unfiltered number of paragraphs. "%used" is "#src" divided by "total".
Appendix B Example of an untranslatable word
Catford (1965) explained the definition of an untranslatable word with an example from French literature La Chatte (The Cat) by Colette:
Fr: “Le soleil allume un crépitement d’oiseaux dans les jardin”
En 1: “The sun kindles a crackling of birds in the garden”
En 2: “the sun... As it rose, already burning hot, it awoke a twittering of sparrows in the gardens”
In the French sentence, the word crépitement is translated as crackling in English by the first translator, but it does not capture the association of crépitement (crackling) with pépiement (twittering) due to shared lexical component in French. The second English translation, though lengthier, captures both the direct meaning and the associated concept of the word “crépitement”. To be aware of this artful play of words, one needs to have in-depth knowledge of the French language.
Appendix C Example source texts and target texts from Par3 dataset
source language source text human references
French Le grand chien noir, qui était ref_1 The big black dog, who had gone over to
allé flairer la bière, revint sniff at the bier, came back reluctantly.
en rechignant. ref_2 The big black dog, who had gone to sniff at
had gone to sniff at the coffin, came back sulkily.
Czech Kadet Biegler viděl, že je ref_1 Cadet Biegler, seeing that the conversation
rozmluva s ním skončena, was at an end, saluted and, very red in the face,
zasalutoval a celý červený passed through the carriage to the corridor at the
v obličeji prošel vagónem, very end.
až se ocitl na samém konci ref_2 Cadet Biegler saw that the conversation
v příčné chodbě. with him was over, so he saluted and all red in the
face walked through the car until he found himself
at the very end, in the crosswise passageway.
Swedish Ty det var just nu det borde ref_1 For now is when it should have happened.
ha skett. Den som vill Someone who wants to act must be able to seize
handla måste kunna gripa the moment. Who knows if it will return? I’m not
tillfället. Ingen vet om det up to it!
kommer igen. Jag duger ref_2 For it was just now it should have happened.
inte! He who wants to act must seize his chance. No
one knows whether it will ever come back. I’m
not up to it!
Table 5: Examples of source texts and human references from the Medium All dataset where there are meaningful variations between human references for each source text.