### v2

# Multiple References with Meaningful Variations Improve Literary Machine Translation

Abstract
While a source sentence can be translated in many ways, most machine translation (MT) models are trained with only a single reference. Previous work has shown that using synthetic paraphrases can improve MT. This paper investigates best practices for employing multiple references by analyzing the semantic similarity among different English translations of world literature in the Par3 dataset (Thai et al., 2022). We classify the semantic similarity between paraphrases into three levels: low, medium, and high, and fine-tune three different models (mT5-large, LLaMA-2-7B, and Opus-MT) for literary MT tasks. Across different models, holding the total training instances constant, single-reference but more source texts only marginally outperforms multiple-reference with half of the source texts. Moreover, when fine-tuning an LLM, using paraphrases with medium and high semantic similarity outperforms an unfiltered dataset, with improvements in BLEU (0.3–0.5), COMET (0.1–0.9), and chrF++ (0.17–0.32). Our code is publicly available on GitHub. [^1] https://github.com/swsiwu/multi_ref_literary_MT
Multiple References with Meaningful Variations Improve
Literary Machine Translation
Si Wu Northeastern University siwu@ccs.neu.edu                        John Wieting Google DeepMind jwieting@google.com                        David A. Smith Northeastern University dasmith@ccs.neu.edu
1 Introduction
A sentence can be translated into another language in many ways. The differences among translations can be trivial—replacement of a synonym, punctuation, possessive ’s vs. of—or they can be meaningful and reveal different aspects of the source language text. (See examples in Figure 1, Table 6, and Appendix B.)
Literary translators have to consider multiple aspects of a source language when translating: meaning, form, rhyme, style, corresponding culture that the language is embedded in, etc. (Lefevere, 1992). Some emphasize lexical and syntactic imitation (formal equivalence), while others sacrifice form for comprehensibility (dynamic/functional equivalence) (Nida, 1964). Read together, different translations’ emphases and interpretations provide a more comprehensive impression of the source text and a deeper understanding of the source language.
Figure 1: An example of meaningful variations between different English translations of a Classical Chinese source text. Highlighted texts are segments where the variations are meaningful. The example source text and its English translations are from the Par3 dataset (Thai et al., 2022).
Previous work has shown that using synthetic paraphrases can improve MT performance (Madnani et al., 2008; Zheng et al., 2018; Khayrallah et al., 2020), but they do not explore different semantic similarity of paraphrases and their impact on MT performance.
To systematically investigate this, we use the Par3 dataset, which has paragraph-level, complex English paraphrases from literature translated by expert translators (Thai et al., 2022). While there are highly similar paraphrases in Par3, many variations between paraphrases in Par3 are more than replacement, insertion, deletion, and re-ordering of words. These naturally existing paraphrases are more complex than what’s generated by traditional paraphrasers that are rule-based (McKeown, 1983) and SMT-based (Wubben et al., 2010; Bannard and Callison-Burch, 2005; Madnani et al., 2007). While LLMs have improved synthetic paraphrase’s quality and complexity, they are still at risk of mistakes, especially for outputs longer than one sentence (Jayawardena and Yapa, 2024a, b; Su et al., 2022; Horvitz et al., 2024; Mishra et al., 2024). Moreover, non-expert translations are prone to translationese, where the translation is overly literal and awkward (NLLB Team et al., 2022).
In this paper, we investigate two questions: when holding the number of fine-tuning data constant, is using multiple references better than just one target reference? What kind of paraphrase should be incorporated to achieve the best performance?
To address these questions, we construct datasets that vary in the number of source texts, the number of references per source text (single vs. multiple references), and, for multiple-reference datasets, in the level of semantic similarity among the references. We use Paragram-SP (Wieting et al., 2022), an efficient and effective model for measuring semantic similarity. We categorize paraphrases into three groups based on their semantic similarity: low, medium, and high, where medium has the most interesting and meaningful variations among the references for a source text. We then fine-tune three different models: mT5-large, LLaMA-2-7B, and Opus-MT, using the aforementioned datasets on MT tasks and compare their performances against baseline and commercial models: No Language Left Behind (NLLB Team et al., 2022) and Google Translate.
Our findings in the literary domain are:
  • Holding the total training instances constant, single-reference but more source texts only marginally outperforms multiple-reference. (Section 6.1)
  • A dataset of source texts with medium and high semantic similarity among references can outperform an unfiltered dataset even when the unfiltered dataset has more training instances. (Section 6.2)
  • In contrast, training with source texts that have medium and low semantic similarity among references results in worse performance than an unfiltered dataset. (Section 6.2)
The remainder of our paper is organized as follows: we describe our method for measuring semantic similarity in Section 3, then we describe how we construct different datasets based on their semantic similarity and number of source and reference texts in Section 4. We provide details on the models we use for fine-tuning as well as the baseline MT models in Section 5, where we also provide information about the MT metrics we use and relevant training details. Finally, we discuss our experimental results in Section 6.
2 Related work
Paraphrases have been used to augment training data and improve MT performance. Madnani et al. (2008) explore how much improvement one can get by increasing the number of synthetic references when tuning an SMT system. The paraphrases they use are automatically generated using phrase table pivoting. Similarly, Callison-Burch et al. (2006) also use pivot-based method to generate paraphrases to improve SMT performance. However, neither of these works examines the linguistic variations between the reference translations of a source text, which is reasonable since the references are synthetic and much less complex.
Khayrallah et al. (2020) use simulated paraphrases to improve MT performance on low-resource languages. They focus on the general effect of augmenting training data using synthetic paraphrases when training data is scarce. Our work further studies the semantic similarity of paraphrases on MT. It is worth noting that the languages present in our dataset are not low-resource languages, and instead of training a language model from scratch, we fine-tune three different types of models: mT5 (a text-to-text encoder-decoder model), LLaMA-2-7B (an autoregressive language model), and Opus-MT (a bidirectional transformer model for machine translation).
Zheng et al. (2018) propose a method to compress existing human references into a lattice and then generate synthetic references for training to boost MT and text generation performance. They also explore different ways of utilizing multiple references during training with a bidirectional language model. Instead of exploring methods of creating synthetic references, our work uses all available expert references for each source text.
Many studies have operationalized paraphrase taxonomy with different levels of granularity (Bhagat and Hovy, 2013; Wahle et al., 2023; Wegmann et al., 2024). Most are trained on human-annotated paraphrase datasets with explicit definitions and follow different agreement protocols among human annotators. Wegmann et al. (2024) operationalize context-dependent paraphrases with plausible variations in dialog, and they demonstrate their hands-on training approach for annotators. With this annotated dataset, they use in-context learning to train a classifier to detect paraphrases in dialog. In our paper, instead of using human annotations, we use Paragram-SP, a semantic textual similarity scorer, to classify the type of paraphrase, as we are interested in a computational, low-cost, yet effective approach.
Others explore different factors that can impact translation quality: Khayrallah and Koehn (2018) explore how different kinds of noise in training data impact MT performance when using an NMT model. Maillard et al. (2023) show that the addition of high-quality data can significantly improve MT performance on low-resource languages, even when the quantity of high-quality training data is small. Consequently, they emphasize the value of collecting high-quality data from native speakers for training MT models.
3 Measuring semantic similarity
We want to investigate what level of semantic similarity among the references (of a source text) is best for training MT systems by first measuring and classifying it.
Given a source text, we use a semantic textual similarity scoring model, Paragram-SP (Wieting et al., 2022), to measure the semantic similarity between its English translations. Paragram-SP score ranges from -1 to 1, where a lower score indicates low semantic similarity, and a higher score indicates high semantic similarity. A model similar to this one from Wieting et al. (2019b) was successfully used as a reward function in minimum risk training Shen et al. (2016) to improve machine translation quality Wieting et al. (2019a).
In the literary translation dataset Par3 (Thai et al., 2022), each source text has at least two different English translations. Paragram-SP can only score the semantic textual similarity for two texts, so in the case where there are more than two translations, we take the average pairwise Paragram-SP score as the source text’s sim_p score. Mathematically:
[…]
where $M$ is the set of human translations given a source text $s$, and $C$ is all possible combinations of a set of pairs in $M$. Paragram-SP (a,b) is the Paragram-SP score given a pair of English translations $a$ and $b$.
Figure 2: The distribution of sim_p scores for the source texts in the Par3 corpus. Paraphrases with higher scores exhibit greater semantic and syntactic similarity. Upon manual inspection, we observe that the paraphrases within the red and blue vertical lines have meaningful variations: different enough to be interesting, yet similar enough to retain their nature as paraphrases.
To identify the range for meaningful variations, we rank the source texts by their sim_p scores from low to high with their target references attached. Upon manual inspection, we define the range for meaningful variations to be [0.45-0.85], and we call this set of paraphrases the Medium dataset. Subsequently, the lower range [-1,0.45] is the Low dataset, and [0.85,1] is the High dataset.
We provide a detailed description of each sim_p range in Table 1, and the distribution of sim_p in the entire unfiltered dataset is in Figure 2.
sim_p Description Dataset
$1.0$ Completely identical. Medium+High
Unfiltered
$[0.9-1.0)$ Almost identical, except for minor punctuation differences, Medium+High
e.g., single vs double quotation marks, or word replacement Unfiltered
(by a synonym or a phrase).
$[0.85-0.9)$ Differences happen more often, but they are still variations Medium+High
that are not meaningful. Compared to the 0.9-1.0 similarity, Unfiltered
the difference can be clausal instead of just one or two words.
However, the overall amount of information is about the same.
$[0.45-0.85)$ Very meaningful variations. The paraphrases are more likely Medium
to be visually dissimilar. Sometimes, the amount of information Medium+High
can differ much more: some translators are more verbose, Medium+Low
others are more succinct. Unfiltered
$[-1-0.45)$ Noisy paraphrases that have an overall information mismatch. Medium+Low
Paraphrases can be misaligned or erroneous. Unfiltered
Table 1: Different levels of semantic similarity score (sim_p) and their characteristics. See Section 3 for how we compute sim_p given a source text and its reference translations. The dataset column indicates the datasets that include paraphrases from the corresponding sim_p range.
4 Datasets
We classify reference translations based on their semantic similarity with the method described in Section 3, then we create different datasets composed of various sim_p for experiments.
To systematically study the impact of the number of source texts, the number of references, and the type of paraphrase included, we create different training datasets such that each varies one potentially influential factor.
We use #source to denote the number of unique source texts, and #instance denotes the total number of training instances, e.g., 3 source texts each has 3 reference translations, will be denoted as #source = 3, and #instance = 9.
We create four different types of datasets:
  • Single: Each source text has one reference.
  • Medium: Each source text has multiple references, where the semantic similarity between references is medium.
  • Unfiltered: Each source text has multiple references, with semantic similarity between references varying across the dataset.
  • Medium+Low, Medium+High: Each source text has multiple references. These datasets are constructed by starting with the entire Medium dataset and then incrementally adding 10% of Low or High similarity references until references are exhausted.
For each dataset type above except the last type, we create subsets with different numbers of sources or instances by performing random sampling within each dataset category.
The source texts and human references are from Par3 (Thai et al., 2022). [^2] Par3 dataset can be downloaded from https://github.com/katherinethai/par3 The source languages include Portuguese, Swedish, Italian, Czech, Japanese, Classical Chinese, Norwegian, German, Russian, and French. All references are only in English; In other words, the language pairs in this dataset are to-English language pairs. Details on the composition of the Medium dataset compared to the Unfiltered dataset are provided in Appendix 5.
We use paragraphs as units for source and target and only select source paragraphs whose human translations are all under the 128-word limit. Voigt and Jurafsky (2012) have shown that when applying MT to literary translation, using above sentence-level translation can improve literary cohesion. Moreover, a paragraph provides enough length for interesting variations but also increases the complexity of the task.
To ensure train/val/test has the same language split, we split data into train and val/test in 0.8/0.2 ratio for each language by the number of source paragraphs instead of the source-target pairs. Train and val/test also don’t share the same books. Val and test datasets are created by first randomizing at the source paragraph level. The test dataset is used in all experiments, and it has multiple human references per source text for evaluation. It is composed of medium semantic similarity only.
5 Experiments
We investigate the impact of varying the number of source texts, references, and semantic similarity of reference translations on MT tasks in the literary domain. Using datasets described in Section 4, we fine-tune three different models for translation: mT5-large, LLaMA-2-7B, and Opus-MT.
To calibrate our interpretation of the performance of these models, we compare them to out-of-the-box translation results from No Language Left Behind (NLLB) and Google Translate (§5.1). We employ three different translation evaluation metrics (§5.2). In the remainder of this section, we describe our training protocols (§5.3).
5.1 Models
mT5-large
: An encoder-decoder model that covers 101 languages (Xue et al., 2021) including all languages in Par3. It has about 1.2 billion parameters. While not specifically trained for translation, it can be fine-tuned for multilingual translation and other downstream tasks. We fine-tuned this model using Hugging Face Trainer class. [^3] https://huggingface.co/google/mt5-large
LLaMA-2-7B
: An English-centric autoregressive transformer model with limited multilingual capabilities (Touvron et al., 2023). Studies have shown that LLaMA-2 can translate high-resource non-English language with some fine-tuning due to the presence of the corresponding language in the training dataset. LLaMA-2 is also able to adapt to a new low-resource language with some fine-tuning. All the languages in this paper are in the LLaMA-2 training data for >= 0.05%, except that the Chinese in Par3 is classical Chinese, which shares mostly the same characters as modern Chinese but has different grammar. It would be interesting to observe the translation performance after fine-tuning on our datasets given that non-English data is only a minuscule part of LLaMA-2 training data.
Opus-MT
: An open-source multilingual machine translation model. The specific version we use is opus-mt-mul-en, which supports multiple source languages and English as the only target language. All models that are part of the Opus-MT project (Tiedemann et al., 2023; Tiedemann and Thottingal, 2020) are trained on the OPUS corpus using the Marian framework (Junczys-Dowmunt et al., 2018), with an architecture similar to BART (Lewis et al., 2020). Unlike mT5-large and LLaMA-2-7B, Opus-MT is not a large language model, and it performs worse than the previous two. Similar to mT5, it’s also encoder-decoder architecture, but it leverages both bidirectional and autoregressive training objectives. This model is available on Hugging Face. [^4] https://huggingface.co/Helsinki-NLP/opus-mt-mul-en
No Language Left Behind (NLLB):
An open-source MT system that can translate any language pairs between 202 languages, including low-resource languages, such as Asturian, Luganda, Urdu (NLLB Team et al., 2022). All the NLLB models are able to translate out-of-the-box, unlike mT5 which requires fine-tuning for MT downstream tasks. To compare to a fine-tuned mT5-large (1.2B), we choose the NLLB-200-Distilled model with 1.3B parameters. [^5] https://huggingface.co/facebook/nllb-200-distilled-1.3B We choose NLLB as a baseline model to compare to.
Google Translate:
A closed-sourced commercial system. As far as we know, it is trained on data from the public web with a transformer encoder and an RNN decoder. [^6] https://research.google/blog/recent-advances-in-google-translate/ Google Translate’s output is provided in the Par3 dataset.
5.2 Evaluation metrics
BLEU
(Papineni et al., 2002): An n-gram-based MT evaluation metric. Notice that we are using multiple human references for each source input.
$\text{COMET}_{22}^{\text{DA}}$
(Rei et al., 2020): An MT evaluation metric that is based on sentence embeddings. We use wmt22-comet-da by Unbabel (Rei et al., 2022). [^7] https://huggingface.co/Unbabel/wmt22-comet-da It was trained on a transformer-based multilingual masked language model, XLM-R (Conneau et al., 2020). Since $\text{COMET}_{22}^{\text{DA}}$ only accepts one reference per source text for evaluation, for each source text, we take the average of their $\text{COMET}_{22}^{\text{DA}}$ scores regarding different references translations.
chrF++
(Popović, 2017): An MT evaluation metric that is based on character and word-level n-gram recall and precision. We include this metric as Zouhar et al. (2024) find that COMET can struggle in domains outside WMT, e.g., literary domain.
#SRC #INST BLEU $\text{COMET}_{22}^{\text{DA}}$ chrF++
NLLB-1.3B 27.3 68.9 41.62
Google Translate 35.8 72.6 48.02
mT5 LLaMA Opus mT5 LLaMA Opus mT5 LLaMA Opus
single 50k 50k 29.9 30.1 23.2 71.5 70.3 66.6 43.44 42.08 38.28
medium 21k 50k 29.8 30.0 23.0 71.3 72.8 65.9 43.20 44.06 37.99
single 58k 58k 30.4 30.2 23.9 71.5 72.1 67.0 43.81 42.86 38.63
medium 24k 58k 29.8 30.2 23.1 71.3 72.4 66.0 43.36 43.79 38.06
single 77k 77k 31.3 29.4 24.3 71.8 71.3 67.1 44.33 42.81 38.97
medium 32k 77k 30.6 28.2 23.7 71.6 70.2 66.5 43.73 41.12 38.50
Table 2: To assess the efficiency of using multiple references, we keep the total number of training instances (#instance) constant. We compare the performance between each Single (where each source text has only one single human reference) and Medium (where each source text has more than two human references). A paired bootstrap resampling statistical significance test is performed on each pair, with the Single in each pair serving as the baseline (Koehn, 2004). All results are statistically significant ($p<0.01$). Across different models and metrics, Single is only marginally better than Medium, despite having at least twice as many source texts (#source).
#SRC #INST BLEU $\text{COMET}_{22}^{\text{DA}}$ chrF++
NLLB-1.3B 27.3 68.9 41.62
Google Translate 35.8 72.6 48.02
mT5 LLaMA Opus mT5 LLaMA Opus mT5 LLaMA Opus
unfiltered 77k 181k 31.9 30.8 25.3 71.9 72.1 67.5 44.67 44.11 39.54
medium 50k 118k 31.5 30.3 24.5 71.9 71.6 67.1 43.81 43.76 38.95
medium + low 58k 138k 31.1 29.4 24.5 71.7 71.6 67.0 44.16 41.29 38.97
medium + low 59k 140k 31.1 30.1 24.5 71.7 71.7 67.1 44.11 43.64 39.01
medium + high 58k 137k 32.0 29.2 24.7 72.0 70.3 67.4 44.77 42.39 39.21
medium + high 66k 156k 32.2 30.8 25.3 72.1 72.6 67.6 44.92 44.43 39.55
medium + high 68k 159k 32.4 31.1 25.1 72.1 73.0 67.5 44.84 44.31 39.47
Table 3: When working with a dataset where each source text has multiple references, using just Medium (where references have medium semantic similarity) can already yield results comparable to those of the unfiltered dataset. To assess the effect of adding low and high similarity references, we incrementally increase their proportion by 10% until exhausted. The results indicate that using all Medium + High references likely gives the best performance. We tested statistical significance using paired bootstrap resampling, with the unfiltered dataset as the baseline. All results are statistically significant ($p<0.01$) unless marked with $\dagger$.
5.3 Training details
Given many mT5 and LLaMA already have strong language abilities and only require fine-tuning for MT downstream tasks, we consider fine-tuning as a more efficient approach than training from scratch, especially since our dataset is English-centric (all reference translations are in English) and these LLMs usually perform better when translating from non-English language into English.
For fine-tuning, each training instance is a source text and one of its human translations, and the training and validation datasets are randolmized, e.g., (source text A, target human translation $A_{1}$), (source text B, target human translation $B_{3}$), (source text C, target human translation $C_{1}$), etc.
We fine-tuned mT5-large with Hugging Face libraries. Our optimal hyperparameters for fine-tuning are $lr=0.00025$, global batch size $=128$, and batch size $=16$. Training took 16 to 48 hours on 4 GPUs depending on the size of the dataset.
For LLaMA-2-7B, we used LoRA-tuning (Hu et al., 2022) with Meta’s LLaMA-recipe. [^8] https://llama.meta.com/llama-downloads/, [^9] https://github.com/meta-llama/llama-recipes We structured each training instance as "source_text ###> target text" following the fine-tuning method from Zhang et al. (2023). Depending on the dataset size, training took 8 to 24 hours.
For Opus-MT, it already has translation ability. We fine-tuned the model using the Hugging Face Trainer class for it to adopt to the literary domain. For each experiment, we fine-tuned the model for 10 epochs with a batch size of 32 on a V100-SXM2 GPU. Depending on the dataset size, fine-tuning took between 1 and 3 hours.
For all models, we use BLEU on the validation set to select the best epoch. Unlike mT5, we find that LLaMA-2’s performance is significantly influenced by fine-tuning hyperparameters (±$\pm 5$ BLEU). Additionally, we observe that increasing the global batch size improves performance as the training data grows. We report the best performance achieved with three global batch sizes (8, 16, 32) for LLaMA-2-7B.
6 Results and Discussion
Using the datasets, models, and MT metrics described in previous sections, we analyze the effectiveness of the number and semantic similarity of reference translations.
6.1 Single reference vs multiple references
Simply adding additional references to a single-reference training set will, unsurprisingly, improve performance, just as will happen with increases in same-quality training data (Madnani et al., 2007, 2008). The more interesting question, however, is when the number of training instances are the same, does it make a difference whether we increase the amount of source text and pair each with one translation, or we have less source text but each source text is paired with multiple translations?
This is not, to our knowledge, an idea that has been previously explored. In Table 2, we compare Medium vs Single under different total number of instances. When #instance is the same, Single on average only marginally outperforms Medium across different models and metrics: ranging 0.0–0.7 in BLEU, (-)2.5–(+)1.1 in COMET, and (-)1.98–(+)1.69 in chrF++. It is surprising considering Medium has less than half of Single’s #source. For each language model, we tested the statistical significance (p < 0.01) of the result by comparing the Medium and Single datasets with the same #instance, using Single as the baseline. The statistical significance was assessed using paired bootstrap resampling method (Koehn, 2004). [^10] For BLEU and chrF++: https://github.com/mjpost/sacrebleu,  [^11] For COMET: https://github.com/Unbabel/COMET
Our results suggest that, with limited source and target texts, increasing the number of references per source text can improve performance similarly to increasing the number of source-target pairs. However, it remains unclear whether this holds when references are synthetically generated by paraphrasers rather than high-quality human references in the literary domain.
6.2 Unfiltered vs. filtered datasets
Despite having more training instances, Unfiltered generally performs worse than Medium+High (#source $=66k,68k$) across models and metrics, with LLMs showing more notable improvements than the NMT model (Table 3).
We test the statistical significance of our result (Koehn, 2004) and find that Med+High (#source $=66k,68k$) outperforms Unfiltered across different MT metrics and models (p-value < 0.01). Specifically, with mT5 and LLaMA, Medium+High (#source $=68k$) outperforms Unfiltered by 0.3–0.5 BLEU, 0.2–0.9 COMET, and 0.17–0.2 chrF++ (Medium+High #source $=66k$ by 0.25–0.32 in chrF++). This implies that when fine-tuning an MT model with multiple references, it is beneficial to compute the sim_p and filter out the ones with low semantic similarity to achieve the optimal MT performance.
Furthermore, we observe that Medium+Low perform worse than Unfiltered across different models and metrics (Table 3). Too much divergence among references of a source text may create confusion and hinder the performance.
BLEU $\text{COMET}_{22}^{\text{DA}}$ chrF++
mT5 LLaMA Opus mT5 LLaMA Opus mT5 LLaMA Opus
medium + 1.0 31.7 28.9 24.6 72.0 71.3 67.3 44.59 42.46 39.16
medium + 0.9-1.0 31.7 27.8 24.8 71.9 70.3 67.3 44.55 40.75 39.29
medium + 0.85-0.9 31.7 31.0 24.9 71.9 72.9 67.4 44.52 44.70 39.30
Table 4: Ablation study on fine-tuning with the highly similar paraphrases (§6.5). 1.0 is pure repetition, 0.9-1.0 is the paraphrases with trivial differences, and 0.85–0.9 is non-trivial but non-meaningful paraphrases. All experiments have the same number of source texts, #source = 58k. All results are statistically significant ($p<0.01$) unless marked with $\dagger$.
6.3 Performance difference between models
The three models we fine-tune for literary MT task differ in architecture: mT5-large and LLaMA-2-7B are LLMs, while Opus-MT is a much smaller NMT model. As expected, Opus-MT’s performance is generally worse than the LLMs. More interestingly, under the same dataset type, mT5-large and Opus-MT performance improves when #instance increases, whereas LLaMA-2-7B is more unstable. For example, Medium’s BLEU, COMET, and chrF++ scores do not improve as we increase #instance under LLaMA-2-7B.
One potential reason for this discrepancy is that, unlike mT5-large and Opus-MT, which are both multilingual models, LLaMA-2-7B as a primary monolingual model. Its MT ability is still not well understood (Xu et al., 2024). On the other hand, as we mentioned in Section 5.3, LLaMA-2-7B is sensitive to the batch size using LoRA-tuning, which can also account for the less stable performance.
6.4 Performance difference between languages
We compare the performance gain between the Unfiltered and Medium+High (#source = 68k) datasets across different languages and metrics, as shown in Figure 3 in the Appendix. Additionally, we evaluate this gain in relation to the amount of fine-tuning data in the Unfiltered dataset.
While among different metrics, Medium+High outperforms Unfiltered (Table 3), under the same model, each language’s metric gain is different. Surprisingly, for all models, more fine-tuning data for a language (e.g., French and Russian have more fine-tuning data) does not result in proportionally more gain.
The gain for languages with less fine-tuning data also seems inconsistent between different models in both the amount and direction, e.g., Classical Chinese mostly has negative gain for mT5-large while much bigger positive gain using LLaMA-2-7B and Opus-MT. However, overall, between metrics, the direction of gain for a language under the same model, is consistent.
6.5 Ablation study on the highly similar paraphrases
The Medium+High datasets in Table 3 are only different by the quantity of high semantic similarity paraphrases. In this ablation study, we further examine the characteristics of these high semantic similarity paraphrases and their impact on MT performance. We classify these high semantic similarity paraphrases into three categories:
  • sim_p = 1.0: completely identical paraphrases (absolute repetition)
  • sim_p [0.91.0$\in$[0.9-1.0)∈ [ 0.9 - 1.0 ): little variation, mostly lexical: 1 or 2 word replacements, deletion, addition, reordering, and punctuation change.
  • sim_p [0.850.9$\in$[0.85-0.9)∈ [ 0.85 - 0.9 ): slightly less trivial variations than the above. More syntactic variations, but still not meaningful.
For each category, we combine the full Medium dataset (#source = 50k) with 10% additional paraphrases from the same category to form a new dataset (#source = 58k). However, the first category (sim_p = 1.0) does not have enough additional instances. To counter that, since the paraphrases in the first category are pure repetition, we create more training instances by randomly selecting source texts from the other two categories and repeating one of its reference translations.
Across different metrics, Medium+0.85–0.9 demonstrates a marginally improved performance with the caveat that under mT5, performance across these different datasets is similar. We tested the statistical significance (p-value < 0.01) of the results in Table 4 with Medium+1.0 as the baseline.
7 Conclusion
Unlike more utilitarian domains of translation, literary works encourages multiple translators to try their hand at improving on their predecessors. To make the most of this valuable resource for training MT systems, we explore the optimal way to employ multiple references when fine-tuning different models for machine translation.
We find using source texts with medium and high semantic similarity among their references outperform an unfiltered dataset and provide the best performance.
Additionally, we observe that when holding the total number of training instances constant, single-reference (having more source texts with one target text each) is only marginally better than multiple-reference (less source text with multiple target texts each). Future research could test these findings in other MT domains.
8 Limitations
  1. 1.
    We chose to fine-tune LLMs as a more efficient approach, given that these LLMs already have strong language ability and only require fine-tuning for downstream tasks. It is unclear if our claims still hold when one trains an LLM from scratch. Informed by the finding by Zhang et al. (2023), we did not use few-shot and zero-shot methods with LLaMA-2 because LoRA-tuning provides the best MT performance, and we considered LoRA-tuning to be the most appropriate to compare with fine-tuning the mT5-large model. Future research can study the performance difference among these approaches.
  2. 2.
    Fine-tuning mT5-large and LLaMA-2-7B usually requires multiple GPUs. Specifically, we trained mT5-large with 4 A100 GPUs since the model is very large and requires at least 30GB RAM. This could make it difficult for other researchers to replicate our experiments.
9 Potential Risks
MT systems are prone to gender-biased translation errors (Stanovsky et al., 2019), which could happen at test time in this work. Our dataset is also very English-centric since Par3 dataset only contains English paraphrases as target texts. English as a high-resource language is more likely to have multiple translations of the same book than low-resource languages, so working on English only will further widen the performance gap between high and low-resource languages; however, we hope our findings on the characteristics of paraphrases can be used in future research to improve MT on low-resource languages.
Appendix A Medium dataset statistics
lang book #src total %used
Portuguese 4 1,021 3,261 0.31
Swedish 3 1,249 2,591 0.48
Italian 2 1,325 1,912 0.69
Czech 4 1,759 2,886 0.61
Japanese 9 1,779 2,483 0.72
Chinese 7 2,199 3,671 0.60
Norwegian 2 3,011 4,756 0.63
German 16 6,136 8,905 0.69
Russian 27 20,003 34,030 0.59
French 32 32,353 51,052 0.63
Table 5: Composition of the Medium (#source $=50k$) dataset. "#src" is the total number of unique source paragraphs selected, while column "total" is the all unfiltered number of paragraphs. "%used" is "#src" divided by "total".
Appendix B Example of an untranslatable word
Catford (1965) explained the definition of an untranslatable word with an example from French literature La Chatte (The Cat) by Colette:
Fr: “Le soleil allume un crépitement d’oiseaux dans les jardin”
En 1: “The sun kindles a crackling of birds in the garden”
En 2: “the sun... As it rose, already burning hot, it awoke a twittering of sparrows in the gardens”
In the French sentence, the word crépitement is translated as crackling in English by the first translator, but it does not capture the association of crépitement (crackling) with pépiement (twittering) due to shared lexical component in French. The second English translation, though lengthier, captures both the direct meaning and the associated concept of the word “crépitement”. To be aware of this artful play of words, one needs to have in-depth knowledge of the French language.
Appendix C Table 6: source texts and target texts from Par3 dataset
source language source text human reference
French Le grand chien noir, qui était ref_1 The big black dog, who had gone over to
allé flairer la bière, revint sniff at the bier, came back reluctantly.
en rechignant. ref_2 The big black dog, who had gone to sniff at
had gone to sniff at the coffin, came back sulkily.
Czech Kadet Biegler viděl, že je ref_1 Cadet Biegler, seeing that the conversation
rozmluva s ním skončena, was at an end, saluted and, very red in the face,
zasalutoval a celý červený passed through the carriage to the corridor at the
v obličeji prošel vagónem, very end.
až se ocitl na samém konci ref_2 Cadet Biegler saw that the conversation
v příčné chodbě. with him was over, so he saluted and all red in the
face walked through the car until he found himself
at the very end, in the crosswise passageway.
Swedish Ty det var just nu det borde ref_1 For now is when it should have happened.
ha skett. Den som vill Someone who wants to act must be able to seize
handla måste kunna gripa the moment. Who knows if it will return? I’m not
tillfället. Ingen vet om det up to it!
kommer igen. Jag duger ref_2 For it was just now it should have happened.
inte! He who wants to act must seize his chance. No
one knows whether it will ever come back. I’m
not up to it!
Table 6: Examples of source texts and human references from the Medium dataset where there are meaningful variations between human references for each source text.
Appendix D Figure 3: performance difference between different languages
Figure 3: Performance difference between different languages under different MT metrics when fine-tuning with mT5-large (1st row), LLaMA-2-7B (2nd row), and Opus-MT (3rd row). Darker dots are from Medium+High (#source $=68k$) dataset, and lighter dots are from Unfiltered dataset. The bar chart shows the number of training instances in Unfiltered dataset for each language.